OpenStudy (anonymous):

Please help for medal- its extra credit. Let f(x) = -5x - 4 and g(x) = 6x - 7. Find f(x) + g(x) A. x - 11 B. -11x - 11 C. x + 3 D. -11x = 3b

OpenStudy (freckles):

(-5+6)x=? -4-7=?

jimthompson5910 (jim_thompson5910):

\[\Large {\color{red}{f(x) = -5x-4}}\] \[\Large {\color{blue}{g(x) = 6x-7}}\] -------------------------------- \[\Large {\color{red}{f(x)}} + {\color{blue}{g(x)}} = ({\color{red}{-5x-4}})+({\color{blue}{6x-7}})\] \[\Large f(x) + g(x) = -5x-4 + 6x - 7\] \[\Large f(x) + g(x) = ???\]

OpenStudy (anonymous):

x-11

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Can you help with another one? I dont understand the whole domain part... Let f(x) = 3x + 2 and g(x) =7x + 6. Find f ·g and its domain. A. 6x2 + 4x + 42; all real numbers B. 21x2 + 32x + 12; all real numbers C. 21x2 +32x + 12; all real numbers except x = -6/7 D. 6x2 + 4x + 42; all real numbers except x = -2/3 I know it's 21x^2+32x +12

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

f * g means f(x) times g(x) so you have to expand out (3x+2)*(7x+6) you can use the FOIL rule or a table like this one